OpenStudy (anonymous):

f(x)=1/x+11 f-1(x)=_______ can someone help me with this? step by step?

OpenStudy (aravindg):

x

OpenStudy (anonymous):

hey can u help me solve that function

OpenStudy (anonymous):

make f(x) y and switch them. y = 1/x+11 <-----> x = 1/y+11 (switch the variable) now solving for y you'll get y inverse... y + 11 = 1/x y = 1/x - 11

OpenStudy (anonymous):

so f^(-1)(x) = (1/x) - 11

OpenStudy (anonymous):

this function says: take the reciprocal, then add 11 inverse says : subtract 11, then take the reciprocal. it is \[f^{-1}(x)=\frac{1}{x-11}\]

OpenStudy (anonymous):

we can do it the other way too, get the same answer. \[x=\frac{1}{y}+11\] \[x-11=\frac{1}{y}\] \[\frac{1}{x-11}=y\] so \[f^{-1}(x)=\frac{1}{x-11}\]

OpenStudy (anonymous):

\[f-1 (X) = 1\div x-11\] i think like thaz...wrong or correct...